# DESI Dark Energy Time Evolution is Recovered by Cosmologically Coupled Black Holes

### v2

@@ Abstract
Recent baryon acoustic oscillation (BAO) measurements by the Dark Energy Spectroscopic Instrument (DESI) provide evidence that dark energy (DE) evolves with time, as parameterized by a $w_{0}w_{a}$ equation of state. Cosmologically coupled black holes (BHs) provide a DE source that naturally evolves with time, because BH production tracks cosmic star-formation. Using DESI BAO measurements and priors informed by Big Bang Nucleosynthesis, we measure the fraction of baryonic density converted into BHs, assuming that all DE is sourced by BH production. We find that the best-fit DE density tracks each DESI best-fit $w_{0}w_{a}$ model within $1\sigma$, except at redshifts $z\lesssim 0.2$, highlighting limitations of the $w_{0}w_{a}$ parameterization. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)~{}\,{\rm km\,s^{-1}\,Mpc^{-1}}$, in excellent agreement with the Chicago-Carnegie Hubble Program, with the same $\chi^{2}$ as $\Lambda$CDM, and with two fewer parameters than $w_{0}w_{a}$. Because cosmologically coupled BH production depletes the baryon density established by primordial nucleosynthesis, these BHs provide a physical explanation for the “missing baryon problem” and the anomalously low sum of neutrino masses preferred by DESI. The global evolution of DE is an orthogonal probe of cosmological coupling, complementing constraints on BH mass-growth from elliptical galaxies, stellar binaries, globular clusters, the LIGO-Virgo-KAGRA merging population, and X-ray binaries. A DE density that correlates with cosmic star-formation: 1) is a natural outcome of cosmological coupling in BH populations; 2) eases tension between early and late-time cosmological probes; and 3) produces time-evolution toward a late-time $\Lambda$CDM cosmology different from Cosmic Microwave Background projections.
@@ 1 Introduction
Note that if $01+w_{0}+w_{a}>$ 0, then the denominator would diverge in the early universe as $a\to 0$. Thus $1+w_{0}+w_{a}\leqslant 0$ must always be true, and the DE density initially increases from zero at $a=0$. The exponential in the numerator then guarantees a “hump” for $w_{a}<0$. Other physical behaviors, like “asymptotes” and “multiple humps,” cannot be described by Eq. 1.1. The DESI data may already highlight the provisional nature of Eq. 1.1, because the preferred values for $w_{0}$ and $w_{a}$ allow causality violation and generate energy densities from nothing. [^1] An equation of state $w<-1$ is often referred to as “phantom” for this reason. Phantom models typically require a negative kinetic energy term, which can be envisioned as a scalar field rolling up its potential [14]. In addition, phantom models have been argued to have theoretical inconsistencies [15, 16, 17, 18]. Because this would imply an unstable vacuum, the interpretation of DESI data within the context of the $\Lambda$CDM model [19, 20, 21, 22, 23, 24], as well as suggestions for interacting DE [25], quintessence scalar fields [26, 27, 28], and other scenarios beyond $\Lambda$CDM [29, 30], has attracted recent interest from the scientific community.
@@ 3 Methods
Our goal is to determine the DESI BAO best-fit background-only CCBH and $\Lambda$CDM cosmologies. By fitting $\Lambda$CDM, we validate our independent pipeline against that of DESI. For CCBH, we fit for the collapse fraction $\Xi$, the free parameter that replaces $H_{0}$ used by DESI when fitting background-only $\Lambda$CDM with a BBN prior. For our primary analysis, we adopt a redshift $z>4$ SFRD that includes contributions from intrinsically faint objects [5, Fig. 3], while accurately reproducing $M_{UV}<-17$ observations from JWST [e.g. 93, Fig. 8] (although [c.f. 94, 95]). For $z\leqslant 4$, we adopt the standard Madau & Dickinson [6, 7] SFRD, enforcing continuity by scaling the normalization to match at $z=4$. The required factor $\sim 2$ is consistent with radio astrometric measures of the SFRD from synchrotron emission in protostellar disks [96]. These measurements are free from dust-obscuration systematics that impact all IR, UV, and optical inference of the SFRD and likely better determine the normalization. This combination of $z>4$ and $z<4$ behavior we will designate as “Trinca $\psi$.” We additionally compute with the standard Madau & Dickinson SFRD, designated as “Madau $\psi$,” to provide a fiducial visual reference in Figure 1.555Empirical determinations of the SFRD necessarily assume a particular cosmology to convert observed luminosities into stellar production rates, and to convert object counts into comoving densities. Because the expansion history does not begin to deviate from $\Lambda$CDM until $z\sim 1$ when star-formation is winding down, and the corrections are small relative to the original data uncertainties, we have neglected this correction.
@@
$r_{\mathrm{d}}=55.154\frac{\exp\left[-72.3\left(\omega_{\nu}^{\mathrm{proj}}+0.0006\right)^{2}\right]}{\left(\omega_{c}^{\mathrm{proj}}+\omega_{b}^{\mathrm{proj}}\right)^{0.25351}\left(\omega_{b}^{\mathrm{proj}}\right)^{0.12807}}~{}\mathrm{Mpc}\,,$ (3.1)
@@ 4 Results
Figure 3 summarizes the posterior distributions from our primary CCBH Trinca $\psi$ and Flat $\Lambda$CDM analyses. For comparison, we also include Planck TT,TE,EE+lowE+lensing posteriors. We have omitted $\omega_{b}^{\mathrm{proj}}$ because the posterior distribution for $\omega_{b}^{\mathrm{proj}}$ tracks the prior (see Table 1). The present-day expansion rate $H_{0}=(69.94\pm 0.81)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ is notably higher than both Flat $\Lambda$CDM (DESI 1YR+BBN) and the Planck value, because the consumption of baryons allows earlier transition to DE domination. We highlight that this value agrees to $0.5\%$ with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported by the Chicago-Carnegie Hubble Program using Cepheids, the Tip of the Red Giant Branch (TRGB), and the J-Region Asymptotic Giant Branch stellar distance-ladder calibrations [107]. In the CCBH Trinca $\psi$ cosmology, the recovered value for $\omega_{c}$ is higher than Planck, allowing this CCBH cosmology to accommodate more background matter density contribution from massive neutrinos. In contrast, the Flat $\Lambda$CDM cosmology requires a lower $\omega_{c}$ than Planck in order to achieve the smaller $\Omega_{m}$ preferred by DESI BAO.
@@ 5 Discussion
In fact, there has been a well-known “missing baryons problem” for over a decade, with measurements indicating a shortfall of about $30\%$ [108]. Our estimated survival fraction is consistent with all contemporary baryon censuses [98, Fig. 1]. This consumption of baryons may help explain other unexpected aspects of the DESI-1YR analysis. DESI-1YR results find a most probable summed neutrino mass $\sum m_{\nu}$ equal to zero [Fig. 11, Left Panel 12](see also [109]), contradicting the computed $\sum m_{\nu}>0.059~{}\mathrm{eV}$ based on $\Delta m_{ij}^{2}$ constraints from neutrino oscillation measurements [e.g. 110, 111]. [^10] This standard constraint follows from algebra $\sqrt{\Delta m_{21}^{2}}+\sqrt{\Delta m_{31}^{2}}=(m_{2}-m_{1})+(m_{3}-m_{1})<m_{2}+m_{3}<\sum m_{\nu},$ (5.1) under the assumption that $m_{\nu}\geqslant 0~{}\forall\nu$ and $m_{i}$ are monotonically increasing in $i$ (normal hierarchy). The summed neutrino mass approaching zero in DESI-1YR can be interpreted as a preference for lower matter density at late times, relative to that inferred from the CMB. [^11] It has also been proposed that a preferred zero summed neutrino mass could arise from new physics in the neutrino sector [112] We suspect that the baryon consumption required in the BH DE scenario will allow $\sum m_{\nu}$ to increase towards physically reasonable values. At background order, $\Omega_{b}h^{2}$ and $\sum m_{\nu}$ are not entirely degenerate in a CCBH cosmology, as the required DE density constrains the amount of baryon consumption. We note that $\sum m_{\nu}$ is also sensitive to the growth function at first-order in cosmological perturbations, which can be altered by novel CCBH dynamics [e.g. 64, §4]. We further expect that the decreased baryon density along the line of sight will slightly increase $z_{\mathrm{reio}}$, because $\tau_{\mathrm{reio}}$ is predominantly determined by the low $\ell$ multipoles of the CMB EE polarization power-spectrum. A comprehensive investigation of these, and other first-order observables, is the subject of future work.
@@
To place the measured value of $\Xi$ into context, we assume a Chabrier [113] distribution of stellar masses at birth $\mathrm{d}N/\mathrm{d}m\propto m^{-2.3}$. Then the fraction of stellar mass in stars large enough to form BHs ($m\gtrsim 20M_{\odot}$) is $f=0.4$. A $\Xi=1.4$ implies approximately 3.45×3.45$\times$3.45 × additional baryonic consumption, which is reasonable given astrophysical uncertainties. For example, the stellar Initial Mass Function (IMF) at $z\gtrsim 1$ is an active area of investigation [e.g. 114, 115]. If the IMF at $z\gtrsim 1$ were ‘top-heavy,’ as suggested by some studies [116, 117, 118, 119], the required factor decreases. For example, an IMF $\propto m^{-2.1}$ [e.g. 120] gives $f=0.74$, reducing the factor to 1.86×1.86$\times$1.86 ×. Post-collapse accretion could account for part of $\Xi$, though estimating its extent is challenging. Accretion onto isolated BHs is believed to be inefficient. Because $\Xi$ describes average behavior, a scenario where the high-end of the BH mass function accretes by a factor of $\sim 10^{2}\text{\textendash}10^{3}\times$, while the low end evolves passively, is conceivable. Gravitational processes can also produce mass. This physics is independent of $\Xi$ because Eq. 2.4 has no knowledge of local gravitational processes. The Press and Teukolsky process [121, 122] can convert BH spin into mass, provided the radiated energy is absorbed by further infalling material during core collapse. Local strong gravitational effects would effectively amplify the term on the right-hand side of Eq. 2.4, without altering the depletion described by Eq. 2.2. We note that such effects may involve mediating fields, such as scalar fields [123, 124].
@@
We have focused on a “cosmic noon” scenario, but production of DE at the peak of star-formation complements production near “cosmic dawn” $z\sim 20$, as has been previously studied [125, 63, 64, 68]. The impact of Pop III [e.g. 126] and direct-collapse BHs [e.g. 127] is to establish an effective DE “floor.” Production of this DE floor does not significantly deplete the baryon density, because the DE produced per unit baryon density scales as $(1+z)^{3}$. For example, in Planck 2018 cosmologies, $\omega_{\Lambda}\sim 5\omega_{c}/2$. Assume a cosmic dawn burst of production, either direct collapse or Pop III stellar collapse, at $z_{\mathrm{III}}$. Then
@@
In summary, DESI’s first year of observations provide evidence that DE evolves over time using a $w_{0}w_{a}$ parameterization. If DE is produced from the conversion of baryonic matter into cosmologically coupled BHs, the DE density will increase as massive stars form and collapse into BHs, rising from zero toward a plateau as star-formation quenches at $z\lesssim 2$. We have tested this hypothesis by fitting a suitable Friedmann cosmology to the DESI measured Baryon Acoustic Oscillation signals between $2.50<z<$ 2.5. Our calculated evolving DE density agrees with the DESI $w_{0}w_{a}$ models at $<1\sigma$, except at redshifts $z\lesssim 0.2$, where the $w_{0}w_{a}$ parameterization becomes inadequate for our phenomenology. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)\,{\rm km\,s^{-1}\,Mpc^{-1}}$, in excellent agreement with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported very recently by Chicago-Carnegie Hubble Program using Cepheids, Tip of the Red Giant Branch (TRGB), and J-Region Asymptotic Giant Branch stellar distance-ladder calibrations. They provide a physical explanation for this larger central value: the Hubble rate today is higher than that inferred from the Cosmic Microwave Background because the conversion of baryons into DE during star-formation causes an earlier transition from matter to DE dominance. In addition to reducing tensions between early and late-time observations, the scenario we have investigated provides an astrophysically motivated explanation for a known $\sim 30\%$ deficit in the present-day baryon density, relative to BBN expectations. Furthermore, the conversion of baryons into DE decreases the present-day matter density of Friedmann models, allowing the sum of neutrino masses to increase away from zero. No other current DE model is both astrophysically motivated and capable of reconciling early and late universe measurements of the expansion rate.

### v3

@@ Abstract
Recent baryon acoustic oscillation (BAO) measurements by the Dark Energy Spectroscopic Instrument (DESI) provide evidence that dark energy (DE) evolves with time, as parameterized by a $w_{0}w_{a}$ equation of state. Cosmologically coupled black holes (BHs) provide a DE source that naturally evolves with time, because BH production tracks cosmic star-formation. Using DESI BAO measurements and priors informed by Big Bang Nucleosynthesis, we measure the fraction of baryonic density converted into BHs, assuming that all DE is sourced by BH production. We find that the best-fit DE density tracks each DESI best-fit $w_{0}w_{a}$ model within $1\sigma$, except at redshifts $z\lesssim 0.2$, highlighting limitations of the $w_{0}w_{a}$ parameterization. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)~{}\,{\rm km\,s^{-1}\,Mpc^{-1}}$, with the same $\chi^{2}$ as $\Lambda$CDM, and with two fewer parameters than $w_{0}w_{a}$. This value reduces tension with SH0ES to $2.7\sigma$ and is in excellent agreement with recent measurements from the Chicago-Carnegie Hubble Program. Because cosmologically coupled BH production depletes the baryon density established by primordial nucleosynthesis, these BHs provide a physical explanation for the “missing baryon problem” and the anomalously low sum of neutrino masses preferred by DESI. The global evolution of DE is an orthogonal probe of cosmological coupling, complementing constraints on BH mass-growth from elliptical galaxies, stellar binaries, globular clusters, the LIGO-Virgo-KAGRA merging population, and X-ray binaries. A DE density that correlates with cosmic star-formation: 1) is a natural outcome of cosmological coupling in BH populations; 2) eases tension between early and late-time cosmological probes; and 3) produces time-evolution toward a late-time $\Lambda$CDM cosmology different from Cosmic Microwave Background projections.
@@ 1 Introduction
Note that if $01+w_{0}+w_{a}>$ 0, then the denominator would diverge in the early universe as $a\to 0$.
Thus $1+w_{0}+w_{a}\leqslant 0$ must always be true, and the DE density initially increases from zero at $a=0$. The exponential in the numerator then guarantees a “hump” for $w_{a}<0$. Other physical behaviors, like “asymptotes” and “multiple humps,” cannot be described by Eq. 1.1. The DESI data may already highlight the provisional nature of Eq. 1.1, because the preferred values for $w_{0}$ and $w_{a}$ allow causality violation and generate energy densities from nothing. [^1] An equation of state $w<-1$ is often referred to as “phantom” for this reason. Phantom models typically require a negative kinetic energy term, which can be envisioned as a scalar field rolling up its potential [14]. In addition, phantom models have been argued to have theoretical inconsistencies [15, 16, 17, 18]. Because this would imply an unstable vacuum, the interpretation of DESI data within the context of the $\Lambda$CDM model [19, 20, 21, 22, 23, 24], as well as suggestions for interacting DE [25], quintessence scalar fields [26, 27, 28], and other scenarios beyond $\Lambda$CDM [29, 30], has attracted recent interest from the scientific community.
@@ 3 Methods
Our goal is to determine the DESI BAO best-fit background-only CCBH and $\Lambda$CDM cosmologies. By fitting $\Lambda$CDM, we validate our independent pipeline against that of DESI. For CCBH, we fit for the collapse fraction $\Xi$, the free parameter that replaces $H_{0}$ used by DESI when fitting background-only $\Lambda$CDM with a BBN prior. For our primary analysis, we adopt a redshift $z>4$ SFRD that includes contributions from intrinsically faint objects [11, Fig. 3], while accurately reproducing $M_{UV}<-17$ observations from JWST [e.g. 93, Fig. 8] (although [c.f. 94, 95]). For $z\leqslant 4$, we adopt the standard Madau & Dickinson [12, 13] SFRD, enforcing continuity by scaling the normalization to match at $z=4$. The required factor $\sim 2$ is consistent with radio astrometric measures of the SFRD from synchrotron emission in protostellar disks [96]. These measurements are free from dust-obscuration systematics that impact all IR, UV, and optical inference of the SFRD and likely better determine the normalization. This combination of $z>4$ and $z<4$ behavior we will designate as “Trinca $\psi$.” We additionally compute with the standard Madau & Dickinson SFRD, designated as “Madau $\psi$,” to provide a fiducial visual reference in Figure 1.555Empirical determinations of the SFRD necessarily assume a particular cosmology to convert observed luminosities into stellar production rates, and to convert object counts into comoving densities. Because the expansion history does not begin to deviate from $\Lambda$CDM until $z\sim 1$ when star-formation is winding down, and the corrections are small relative to the original data uncertainties, we have neglected this correction.
@@
$r_{\mathrm{d}}=55.154\frac{\exp\left[-72.3\left(\omega_{\nu}+0.0006\right)^{2}\right]}{\left(\omega_{c}+\omega_{b}^{\mathrm{proj}}\right)^{0.25351}\left(\omega_{b}^{\mathrm{proj}}\right)^{0.12807}}~{}\mathrm{Mpc}\,,$ (3.1)
@@ 4 Results
Figure 3 summarizes the posterior distributions from our primary CCBH Trinca $\psi$ and Flat $\Lambda$CDM analyses. For comparison, we also include Planck TT,TE,EE+lowE+lensing posteriors. We have omitted $\omega_{b}^{\mathrm{proj}}$ because the posterior distribution for $\omega_{b}^{\mathrm{proj}}$ tracks the prior (see Table 1). The present-day expansion rate $H_{0}=(69.94\pm 0.81)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ is notably higher than both Flat $\Lambda$CDM (DESI 1YR+BBN) and the Planck value, because the consumption of baryons allows earlier transition to DE domination. With respect to the latter, CCBH reduces gaussian tension in $H_{0}$ from $5.6\sigma$ to $2.7\sigma$ when compared to the SH0ES measurement [107]. We highlight that this value agrees to $0.5\%$ with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported by the Chicago-Carnegie Hubble Program using Cepheids, the Tip of the Red Giant Branch (TRGB), and the J-Region Asymptotic Giant Branch stellar distance-ladder calibrations [108]. In the CCBH Trinca $\psi$ cosmology, the recovered value for $\omega_{c}$ is higher than Planck, allowing this CCBH cosmology to accommodate more background matter density contribution from massive neutrinos. In contrast, the Flat $\Lambda$CDM cosmology requires a lower $\omega_{c}$ than Planck in order to achieve the smaller $\Omega_{m}$ preferred by DESI BAO.
@@ 5 Discussion
In fact, there has been a well-known “missing baryons problem” for over a decade, with measurements indicating a shortfall of about $30\%$ [109]. Our estimated survival fraction is consistent with all contemporary baryon censuses [98, Fig. 1]. This consumption of baryons may help explain other unexpected aspects of the DESI-1YR analysis. DESI-1YR results find a most probable summed neutrino mass $\sum m_{\nu}$ equal to zero [Fig. 11, Left Panel 8](see also [110]), contradicting the computed $\sum m_{\nu}>0.059~{}\mathrm{eV}$ based on $\Delta m_{ij}^{2}$ constraints from neutrino oscillation measurements [e.g. 111, 112]. [^10] This standard constraint follows from algebra $\sqrt{\Delta m_{21}^{2}}+\sqrt{\Delta m_{31}^{2}}=(m_{2}-m_{1})+(m_{3}-m_{1})<m_{2}+m_{3}<\sum m_{\nu},$ (5.1) under the assumption that $m_{\nu}\geqslant 0~{}\forall\nu$ and $m_{i}$ are monotonically increasing in $i$ (normal hierarchy). The summed neutrino mass approaching zero in DESI-1YR can be interpreted as a preference for lower matter density at late times, relative to that inferred from the CMB. [^11] It has also been proposed that a preferred zero summed neutrino mass could arise from new physics in the neutrino sector [113] We suspect that the baryon consumption required in the BH DE scenario will allow $\sum m_{\nu}$ to increase towards physically reasonable values. At background order, $\Omega_{b}h^{2}$ and $\sum m_{\nu}$ are not entirely degenerate in a CCBH cosmology, as the required DE density constrains the amount of baryon consumption. We note that $\sum m_{\nu}$ is also sensitive to the growth function at first-order in cosmological perturbations, which can be altered by novel CCBH dynamics [e.g. 64, §4]. We further expect that the decreased baryon density along the line of sight will slightly increase $z_{\mathrm{reio}}$, because $\tau_{\mathrm{reio}}$ is predominantly determined by the low $\ell$ multipoles of the CMB EE polarization power-spectrum. A comprehensive investigation of these, and other first-order observables, is the subject of future work.
@@
To place the measured value of $\Xi$ into context, we assume a Chabrier [114] distribution of stellar masses at birth $\mathrm{d}N/\mathrm{d}m\propto m^{-2.3}$. Then the fraction of stellar mass in stars large enough to form BHs ($m\gtrsim 20M_{\odot}$) is $f=0.4$. A $\Xi=1.4$ implies approximately 3.45×3.45$\times$3.45 × additional baryonic consumption, which is reasonable given astrophysical uncertainties. For example, the stellar Initial Mass Function (IMF) at $z\gtrsim 1$ is an active area of investigation [e.g. 115, 116]. If the IMF at $z\gtrsim 1$ were ‘top-heavy,’ as suggested by some studies [117, 118, 119, 120], the required factor decreases. For example, an IMF $\propto m^{-2.1}$ [e.g. 121] gives $f=0.74$, reducing the factor to 1.86×1.86$\times$1.86 ×. Post-collapse accretion could account for part of $\Xi$, though estimating its extent is challenging. Accretion onto isolated BHs is believed to be inefficient. Because $\Xi$ describes average behavior, a scenario where the high-end of the BH mass function accretes by a factor of $\sim 10^{2}\text{\textendash}10^{3}\times$, while the low end evolves passively, is conceivable. Gravitational processes can also produce mass. This physics is independent of $\Xi$ because Eq. 2.4 has no knowledge of local gravitational processes. The Press and Teukolsky process [122, 123] can convert BH spin into mass, provided the radiated energy is absorbed by further infalling material during core collapse. Local strong gravitational effects would effectively amplify the term on the right-hand side of Eq. 2.4, without altering the depletion described by Eq. 2.2. We note that such effects may involve mediating fields, such as scalar fields [124, 125].
@@
We have focused on a “cosmic noon” scenario, but production of DE at the peak of star-formation complements production near “cosmic dawn” $z\sim 20$, as has been previously studied [126, 63, 64, 68]. The impact of Pop III [e.g. 127] and direct-collapse BHs [e.g. 128] is to establish an effective DE “floor.” Production of this DE floor does not significantly deplete the baryon density, because the DE produced per unit baryon density scales as $(1+z)^{3}$. For example, in Planck 2018 cosmologies, $\omega_{\Lambda}\sim 5\omega_{c}/2$. Assume a cosmic dawn burst of production, either direct collapse or Pop III stellar collapse, at $z_{\mathrm{III}}$. Then
@@
In summary, DESI’s first year of observations provide evidence that DE evolves over time using a $w_{0}w_{a}$ parameterization. If DE is produced from the conversion of baryonic matter into cosmologically coupled BHs, the DE density will increase as massive stars form and collapse into BHs, rising from zero toward a plateau as star-formation quenches at $z\lesssim 2$. We have tested this hypothesis by fitting a suitable Friedmann cosmology to the DESI measured Baryon Acoustic Oscillation signals between $2.50<z<$ 2.5. Our calculated evolving DE density agrees with the DESI $w_{0}w_{a}$ models at $<1\sigma$, except at redshifts $z\lesssim 0.2$, where the $w_{0}w_{a}$ parameterization becomes inadequate for our phenomenology. Cosmologically coupled BHs produce $H_{0}=(69.94\pm 0.81)\,{\rm km\,s^{-1}\,Mpc^{-1}}$, reducing tension with SH0ES and in excellent agreement with $H_{0}=(69.59\pm 1.58)\,\,{\rm km\,s^{-1}\,Mpc^{-1}}$ reported recently by Chicago-Carnegie Hubble Program using Cepheids, Tip of the Red Giant Branch (TRGB), and J-Region Asymptotic Giant Branch stellar distance-ladder calibrations. They provide a physical explanation for this larger central value: the Hubble rate today is higher than that inferred from the Cosmic Microwave Background because the conversion of baryons into DE during star-formation causes an earlier transition from matter to DE dominance. In addition to reducing tensions between early and late-time observations, the scenario we have investigated provides an astrophysically motivated explanation for a known $\sim 30\%$ deficit in the present-day baryon density, relative to BBN expectations. Furthermore, the conversion of baryons into DE decreases the present-day matter density of Friedmann models, allowing the sum of neutrino masses to increase away from zero. No other current DE model is both astrophysically motivated and capable of reconciling early and late universe measurements of the expansion rate.